OpenStudy (ana_98):

how do i simplify sqrt of P^7

OpenStudy (ana_98):

sqrt of n/25?

OpenStudy (ana_98):

and sqrt of y^2 / 7

OpenStudy (ana_98):

which part?

OpenStudy (unknownunknown):

\[(P ^{7})^{1/2}\]

OpenStudy (ana_98):

i dont know how to work that out...sorry...

OpenStudy (unknownunknown):

So we can multiply it out, to reach \[P ^{7/2}\]

OpenStudy (ana_98):

yes, that makes sense. its coming back to me now :)

OpenStudy (unknownunknown):

Okay =)

OpenStudy (ana_98):

shoot. so then would that be the answer or do we have to work it out even more?

OpenStudy (unknownunknown):

P^(7/2) looks simplified enough to me.

OpenStudy (ana_98):

ok wow thanks! what i wanna know is how you got the p7 (1/2 )

OpenStudy (anonymous):

how many times does 2 go in to 7?

OpenStudy (anonymous):

because \[\huge p^{\frac{7}{2}}\] is not the answer you are looking for

OpenStudy (ana_98):

ok makes sense now :) thank you

OpenStudy (anonymous):

let me repeat my self that is NOT the answer

OpenStudy (ana_98):

yes i know haha

OpenStudy (anonymous):

\[\sqrt{p^7}=\sqrt{p^6\times p}=\sqrt{p^6}\sqrt{p}=p^3\sqrt{p}\] is the final answer

OpenStudy (anonymous):

more simply realized as "2 goes in to 7 3 times with a remainder of 1, so \(p^3\) comes outside the radial, \(p\) stays in

OpenStudy (ana_98):

oh okay thanks for verifying

OpenStudy (anonymous):

yw

OpenStudy (ana_98):

so what if i had sqrt of n/25?

OpenStudy (ana_98):

@satellite73

OpenStudy (anonymous):

\[\huge \sqrt{\frac{n}{25}}\] like that?q

OpenStudy (ana_98):

yes :) @satellite73

OpenStudy (anonymous):

then since \(\sqrt{25}=5\) you have \[\frac{\sqrt{n}}{5}\]

OpenStudy (ana_98):

thats it? :p

OpenStudy (anonymous):

yup

OpenStudy (ana_98):

cool. last question is sqrt of y^2/7

OpenStudy (anonymous):

\[\huge \sqrt{y^{\frac{2}{7}}}\]?

OpenStudy (ana_98):

yes its y^2 on the top nad 7 on the bottom

OpenStudy (ana_98):

*and

OpenStudy (ana_98):

@satellite73

OpenStudy (ana_98):



OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

\[\sqrt{y^2}=y\]so first step is \[\frac{y}{\sqrt{7}}\]

OpenStudy (anonymous):

then if you have to rationalize the denominator \[\frac{y}{\sqrt{7}}\times \frac{\sqrt7}{\sqrt7}=\frac{y\sqrt7}{7}\]

OpenStudy (ana_98):

ok cool. thats always the confusing part... so would that be the answer? \[y \sqrt{7}/7 \]

OpenStudy (anonymous):

yes

OpenStudy (ana_98):

thanks !

OpenStudy (anonymous):

yw hope it helped

OpenStudy (ana_98):

it did your very helpful :)